The denominator of a fraction is 4 more than the numerator. If 2 is added to both the numerator and the denominator, the resulting fraction has a value of 2/3. Find the original fraction
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

The denominator of a fraction is 4 more than the numerator. If 2 is added to both the numerator and the denominator, the resulting fraction has a value of 2/3. Find the original fraction
----------------------
n/(n+4) to start
=====
(n+2)/(n+6) = 2/3
3(n+2) = 2(n+6)
etc

Answer by greenestamps(13195)



The response from the other tutor outlines one formal algebraic method for solving the problem.

While it is very likely that a formal algebraic solution was wanted, solving a problem like this informally with logical reasoning and mental arithmetic can be valuable brain exercise.

The denominator is 4 more than the numerator; when 2 is added to both, the denominator is still 4 more than the numerator.

The resulting fraction is equal to 2/3. So we need a fraction equivalent to 2/3 in which the denominator is 4 more than the numerator.

In the fraction 2/3, the denominator is 1 more then the numerator; so multiplying numerator and denominator by 4 will give us an equivalent fraction with the denominator 4 more than the numerator.

So the fraction after adding 2 to both numerator and denominator is 8/12; that means the original fraction was 6/10.

ANSWER: 6/10